# Diffusion-Reinforcement Learning Hierarchical Motion Planning in Adversarial Multi-agent Games

### v1

@@ Abstract
Reinforcement Learning- (RL-)based motion planning has recently shown the potential to outperform traditional approaches from autonomous navigation to robot manipulation. In this work, we focus on a motion planning task for an evasive target in a partially observable multi-agent adversarial pursuit-evasion games (PEG). These pursuit-evasion problems are relevant to various applications, such as search and rescue operations and surveillance robots, where robots must effectively plan their actions to gather intelligence or accomplish mission tasks while avoiding detection or capture themselves. We propose a hierarchical architecture that integrates a high-level diffusion model to plan global paths responsive to environment data while a low-level RL algorithm reasons about evasive versus global path-following behavior. Our approach outperforms baselines by 51.2% by leveraging the diffusion model to guide the RL algorithm for more efficient exploration and improves the explanability and predictability. [^1] Code and implementation details are at https://github.com/CORE-Robotics-Lab/Opponent-Modeling
@@ I Introduction
Multi-agent pursuit-evasion games (PEG) [1] are ubiquitous in real-world robotics scenarios where robotic systems could be applied to tracking drug smugglers [2], defensive escorting [3], or robot soccer [4]. In PEG, one team seeks to track its target while the target agent or team employs evasive maneuvers, and research naturally focuses on developing methods for either the pursuers [5] or evaders [6, 7, 8]. While there has been extensive work on developing search policies, including robust tracking frameworks [9], topology map maintenance [10], and multi-hypothesis filters [11], relatively little attention has been paid to evasion. This lack of emphasis provides an opportunity to design a novel evasive policy which is also useful in the real-world scenarios (e.g. business ship evading from pirates).
Traditional approaches to develop evasion policies typically assume full environment observability and are designed to optimize capture time [12] or an ad hoc utility function [13].
@@
More recently, learning-based methods, such as Graph Neural Networks [14], diffusion models [15] or reinforcement learning (RL) methods [16] have been widely used in the pursuit-evasion games and in similar navigation problems [17, 18, 19] (e.g. goal-reaching, path-planning, etc.). These methods often outperform expert policies [20]. Unfortunately, learning the RL policy is usually subject to two key challenges—partial observability and high sample complexity when exploring in large-scale, partially observable environments, as considered in our work.
Researchers propose hierarchical approaches to help exploration by structuring the goal, policies, or rewards. Goal-oriented reinforcement learning (e.g. hindsight experience replay [21]) increases the reward density by assuming the goal has been reached in each exploration trajectory and correcting observation and reward accordingly. Hierarchical reinforcement learning [22] improves the sample efficiency from spatial or temporal abstraction by multiple hierarchically organized RL policies, and hybrid reward structure based methods [23] try to capture the multi-modal distribution of reward where each part of value function can be a low-dimensional representation. However, their applications are either tested in fully observable games [22] or single agent scenarios [24], and sometimes require expert prior knowledge as primitives [25].
In this paper, we propose a novel hierarchical learning framework that utilizes diffusion and RL as high and low level policies respectively. The diffusion model is used as a global path generator to satisfy static map constraints, such as terminal states and obstacle avoidance. Additionally, we employ a low-level RL policy that learns to follow the waypoints while prudently evading captures (see Figure 1). Our key insight is that using the diffusion model can improve the RL performance by constraining its exploration to high-value state-action regions. We find our method is especially useful for the evader to learn escaping behaviors in the large multi-agent, multi-goal partially observable settings — Prisoner Escape and Narco Interdiction Domain (see §III-B). Increasing understanding of potential evader strategies will help future development of countermeasures.
@@
    Our work is the first to create a learning-based framework for evaders in large partially observable environments that learns without expert domain knowledge.
@@
    We propose a novel hierarchical system consisting of a diffusion model as a high level global path planner and a low level RL agent to learn evasive maneuvers, which significantly outperforms all baselines.
@@
    To the best of our knowledge, we pioneer the use of task-oriented costmap construction to enhance explainability, predictability, and flexibility, thereby facilitating a deeper understanding of the evader’s performance.
@@ II-A Pursuit-Evasion Games
A perfect information PEG (i.e., full observability) can be modeled as a differential game and solved using control theory [26, 27, 28, 29]. In comparison, the incomplete information PEG (i.e., partial observability) is a more realistic setting. Researchers commonly use probabilistic models [30, 31] and formulate PEG as optimization problems with the objective of maximizing the payoff [32, 33]. Therefore, recent research also focuses on deriving the optimal policy via RL [34, 32, 35]. Unfortunately, most of RL work focuses on pursuit policies [34, 32, 36]. While Generative Adversarial Imitation Learning has been employed for training an evader policy, this approach requires a dataset showing expert evasive behaviors and a complex curriculum RL scheme [35]. Furthermore, prior work only validates this approach on a small map. In comparison, our method trains evasive RL policy from scratch without any evasion behavior demonstrations and is simpler and more suitable for large, adversarial, and partially observable domains.
@@ II-B Diffusion Models in Path Generating
Diffusion models are a class of generative models that have demonstrated impressive performance in generating sequential data across several applications, such as image synthesis and video generation [37, 38, 39]. Recently, there have been several works employing diffusion models in trajectory generation tasks. Diffuser [40] trains a diffusion model for trajectory generation on offline datasets and plans future trajectories with guided sampling. Diffusion policy extended this work to imitation learning and was able to learn robust policies for various pushing tasks conditioned only on visual inputs [41]. Recent work [15] has extended these approaches for generating trajectories of multiple opponents. There have also been many following works using diffusion models to synthesize trajectories and augment experience buffer for efficient offline RL training [42, 43]. However, these works only function in fully-observable domains, where building a world model is feasible. We instead use the diffusion model as a trajectory planner that allows the reinforcement learning agent to generate diverse plan options in our partially observable pursuit-evasion domain.
II-C Reinforcement Learning in Motion Planning
Recent works have demonstrated the success of RL for motion planning problems in dynamic environments [44, 45, 46, 47, 17]. Many papers discuss how to use RL to avoid obstacles by analyzing exteroceptive sensor information such as RGB cameras [44, 47], LIDARs [46, 47], or laser data converted from images [45]. However, none of these prior works assumes the an adversarial environment, where the “obstacles” are intelligent pursuers, and the map size is as large as in our scenario.
The most similar work from the system design perspective is from [17] which also employs a hierarchical structure in goal-conditioned planning. However, their environment does not incorporate adversaries or multiple goals. Additionally, they require expert datasets for training their RL policy. Dynamic obstacles are considered in [48]; however, these obstacles can only run on a pre-defined segment with a constant velocity, and all the environment objects are observable such that A* search can be employed effectively. In our scenario, we assume the cameras are unobservable to the evader, while the pursuit team can leverage these cameras to detect and track the evader. Therefore, we use a diffusion model to generate diverse global candidate paths and make it possible for RL to do useful exploration about the map.
@@ III-A Partially Observable Markov Decision Process
We model adversarial search and tracking as a Partially Observable Markov Decision Process (POMDP). A POMDP for an agent can be defined by a state $s$, private observation $o$, action $a$, a transition function $\mathcal{T}:s\times a\mapsto s$. At each timestep $t$, an agent receives an observation $o^{t}$, chooses an action $a^{t}\in\mathcal{A}$, and obtains a reward $r^{t}:s\times a_{i}\mapsto\mathbb{R}$. The initial world state is drawn from a prior distribution $\rho$. In our environment (see §III-B), we assume the evader starts from a random position from in a corner region of the map while the pursuit team starts from a random position on the map.
III-B Prisoner Escape and Narco Interdiction Domains
In our evaluation of the proposed algorithm, we adjust two large-scale pursuit-evasion domains from prior work [14] named Prisoner Escape and Narco Interdiction (Figure 2). In both domains, an evader (prisoner or smuggler) is being tracked by a team of search agents in a partially observable environment. We draw three hideouts each episodes from all the candidate hideouts. The game-ending goal of the evader is to reach any hideout on the map (considered as success), or reach the time limit (considered as failure). The heterogeneous pursuit team consists of static (cameras) and dynamic agents (search parties and helicopters) that collaborate to search and track the evader in both domains.
The maximum speed of the evader is always lower than the pursuit agents, making it crucial for the evader to plan evasive maneuvers to successfully reach the hideouts undetected. The observation of the evader includes the current timestep, hideout locations, mountain locations, pursuer states and evader states. However, the camera locations and visibility map are hidden to the evader. These domains increase the complexity of the task and more closely match real-world dynamics compared to other pursuit-evasion games.
@@ III-C Pursuit Team Behavior
We create a heuristic policy for tracking the evader inspired by previous object searching works [49, 50] and the fact that pursuit teams typically follow some routine patterns. The search agents can choose to either 1) go to the last known location of the evader, 2) intercept the evader, or 3) search in the vicinity of the evader’s last known location. We assume the detection range of searching team $d_{s}$ is parameterized by the agent type $\beta_{type}$, visibility $c_{v}$ and evader speed $s_{e}$: $d_{s}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{e}+\eta)$, which means a higher visibility and evader speed will lead to higher detection range. The pursuer policy is designed to effectively utilize the available information and adapt to a wide variety of different scenarios.
@@ IV Method
The evader’s objective is to reach one of the hideouts on the map and evade detection by the pursuit team, using only partial map knowledge and lacking any information about camera locations or models for whether the pursuers are able to detect the evader. RL or diffusion-based motion planning each partially fits our setting: RL can hypothetically learn to evade through trial-and-error. However, RL has difficulties with exploration in the large, adversarial, and multi-goal domains. On the other hand, diffusion-based motion planning can generate diverse plans that reach the goals; however, diffusion models struggle to learn evasive behaviors and avoid the pursuing agents. This weaknesses is due to the diffusion model is trained with a denoising loss objective that is not readily incorporated into RL [51]. Additionally, the partial observability induces high variance in training.
We propose to reap the benefits of both methods by leveraging diffusion models to guide RL exploration during training and generate path candidates at inference. We break down the hierarchy into the following: learning diffusion-based global plans that meet start, terminal, and obstacle constraints from RRT* paths (§IV-A), employing RL local evasive behaviors to avoid detection on route (§IV-B), and path selection from reversely constructed costmap during inference (§IV-C).
IV-A Diffusion Based Global Plan
In this section, we describe how we leverage diffusion models as a generative motion planner to provide a prior for the RL agent to traverse in large, partially observable, and adversarial environments [52]. Since A* cannot generate the optimal path to guide RL exploration on partially observable maps with adversarial agents[53], we utilize a diffusion model to plan waypoints for an approximate path that the agent should follow. By using the diffusion model, our method is sample-time efficient and enables the downstream reinforcement learning algorithm to outperform prior approaches. The underlying insight is to produce diverse options from the diffusion model for global exploration to deal with map uncertainty.
The diffusion path planning should have three key attributes: 1) global constraints are satisfied (start and terminal locations, obstacles etc.); 2) generated plans are diverse and multimodal and 3) the sampling time of the diffusion model is low. We guide the diffusion sampling procedure using constraint guidance [15] to best satisfy the constraints and use a small number of sampling timesteps to keep the sampling time low. This low sampling time allows us to run the diffusion model during the RL training procedure which is difficult for traditional planners (e.g., RRT* and A*). Additionally, the parallel sampling at inference time provides diverse candidates for the agent to use.
@@ 7       $\theta=\theta+\alpha\nabla_{\theta}\mathcal{L}(\theta)$
We outline the diffusion training process in Algorithm 1. We first generate paths with RRT* (Line 1) — a sample-based motion planner that yields varied paths even for the same initial and target states, which enables the diffusion model to similarly generate diverse paths. Then we uniformly downsample the path into sparse waypoints (Line 1) and train a diffusion model based on these waypoints (Line 1-1). This process differs from Caravalho et al. [19], which trains a diffusion model directly from the dense and smoothed paths. Generating a path with a few waypoints is sufficient enough to act as the high level planner as we expect the RL policy to act as the low-level dynamic actor to move between waypoints. Generating fewer waypoints also allows sampling time to be faster, which is crucial for training in an RL loop.
@@ 3             $(\mu^{i},\Sigma^{i})\leftarrow s_{\theta}(\tau^{i}),\Sigma_{\theta}(\tau^{i})$
The sampling process is similar to our previous work [15] summarized in Algorithm 2 as diffusion path generator $\pi^{d}$. We apply global constraints at each denoising step (Line 2) to help generate a valid path from start to the end location.
IV-B RL Based Adversarial Behavior Learning
In addition to our diffusion model considering the global objectives, it is also important to design an algorithm directly interacting with the environment that can not only accomplish the tasks from the high-level planners but also capture the dynamic patterns in the map to achieve the local adversarial detection avoiding goals. Even the heuristics have been built in our previous works, we believe that a data-driven reinforcement learning algorithm is a more convenient way compared with the man-crafted policy and can perform better within a certain distribution.
@@ IV-B RL Based Adversarial Behavior Learning
Therefore we employ an RL algorithm Soft Actor-Critic (SAC) as the low-level policy with the following aims: 1) basically follow the waypoints from diffusion path and 2) learn to adjust its velocities to evade from searching agents. We summarize the RL training process in Algorithm 3 where we augment the observation with the current desired waypoint from the high-level planner (Line 3, 3). The training objective can be formulated as Equation (1) where $\approx$ means the similar low-level policy can be derived from both sides.

### v2

# Diffusion-Reinforcement Learning Hierarchical Motion Planning in Multi-agent Adversarial Games

@@ Abstract
Reinforcement Learning (RL)-based motion planning has recently shown the potential to outperform traditional approaches from autonomous navigation to robot manipulation. In this work, we focus on a motion planning task for an evasive target in a partially observable multi-agent adversarial pursuit-evasion game (PEG). Pursuit-evasion problems are relevant to various applications, such as search and rescue operations and surveillance robots, where robots must effectively plan their actions to gather intelligence or accomplish mission tasks while avoiding detection or capture. We propose a hierarchical architecture that integrates a high-level diffusion model to plan global paths responsive to environment data, while a low-level RL policy reasons about evasive versus global path-following behavior. The benchmark results across different domains and different observability show that our approach outperforms baselines by 77.18% and 47.38% on detection and goal reaching rate, which leads to 51.4% increasing of the performance score on average. Additionally, our method improves interpretability, flexibility and efficiency of the learned policy. [^1] Code and supplementary details are at https://github.com/ChampagneAndfragrance/Diffusion_RL
@@ I Introduction
Multi-agent pursuit-evasion games (PEG) [1] are ubiquitous in real-world robotics scenarios, where robotic systems could be applied to drug smuggler tracking [2], defensive escorting [3], or robot soccer [4]. In PEG, one team seeks to track its target while the target agent or team employs evasive maneuvers, and research naturally focuses on developing methods for either the pursuers [5] or evaders [6, 7, 8]. Although there has been extensive work on developing search policies [9, 10, 11], relatively little attention has been paid to evasion. This lack of emphasis provides an opportunity to design a novel evasive policy that is useful in real-world scenarios, such as business ships evading pirates and evacuating lost people in national parks.
@@
Recently, learning-based methods, such as Graph Neural Networks [12], diffusion models [13] or reinforcement learning (RL) methods [14] have been widely used in the pursuit-evasion games and similar navigation problems [15, 16, 17] (e.g. goal-reaching, path-planning, etc.). These methods often outperform expert policies [18]. However, learning the RL policy is usually subject to two key challenges—partial observability and high sample complexity when exploring large-scale, partially observable environments, as considered in our work.
Researchers propose hierarchical approaches to help exploration by structuring the goal, policies, or rewards. Goal-oriented reinforcement learning (for example, a replay of the hindsight experience [19]) increases the reward density by assuming that the goal has been reached in each exploration trajectory and corrects observation and reward accordingly. Hierarchical RL [20] improves sample efficiency from spatial or temporal abstraction by multiple hierarchically organized RL policies, and hybrid reward structure-based methods [21] try to capture the multimodal distribution of reward where each part of the value function can be a low-dimensional representation. However, their applications are tested in fully observable games [20] or require expert prior knowledge as primitives [22].
In this paper, we propose a novel hierarchical learning framework that combines diffusion and RL as high- and low-level policies, respectively. The diffusion model is used as a global path generator to satisfy static map constraints, such as terminal states and obstacle avoidance. Furthermore, we employ a low-level RL policy that learns to follow the waypoints while prudently evading captures (see Figure 1). Our key insight is that using the diffusion model can improve the RL performance by constraining its exploration to high-value state-action regions and can help downstream tasks by providing diverse options. We find our method is especially useful for the evader to learn escaping behaviors in the large multi-agent, multi-goal partially observable settings — Prisoner Escape and Narco Interdiction Domain (see §III-B).
@@
    We propose a novel hierarchical system consisting of a diffusion model as a high-level global path planner to aid RL exploration and a low-level RL agent to learn evasive maneuvers, which outperforms all baselines.
@@
    We design an algorithm to implicitly infer the cost map that is used to select paths within our hierarchical motion planning framework to evade the searching agents. Multiple ablation studies show the interpretability, flexibility, efficiency, and generalizability of our method.
@@
    Our work is the first to equip evaders with hierarchical motion planning to learn escaping while reaching the navigation goal in large partially observable environments without expert primitives.
@@ II-A Pursuit-Evasion Games
A perfect information PEG (i.e., full observability) can be modeled as a differential game and solved using control theory [23]. In comparison, the incomplete information PEG (i.e., partial observability) is a more realistic setting since agents usually do not have full knowledge of the environment at all time. Researchers commonly use probabilistic models [24, 25] and formulate PEG as optimization problems to maximize the payoff [26, 27]. Therefore, recent research also focuses on obtaining the optimal policy through RL [28]. Unfortunately, most RL works focus on pursuit policies [29, 30]. Some recent works [31, 32] have focused on evasive behaviors evolved from scratch, but these policies do not need to reach explicit navigation goals. In comparison, we derive an evader policy that can not only escape from pursuers but also reach goals in large, adversarial, and partially observable domains.
@@ II-B Diffusion Models in Path Generating
Diffusion models are a class of generative models that have demonstrated impressive performance in generating sequential data in several applications, such as image synthesis and video generation [33, 34, 35]. Recently, there have been several works employing diffusion models in trajectory generation tasks. Diffuser [36] trains a diffusion model for trajectory generation on offline datasets and plans future trajectories with guided sampling. The diffusion policy extends this work to imitation learning and demonstrates the ability to learn robust policies for various pushing tasks conditioned only on visual input [37]. Recent work [13] has extended these approaches for generating trajectories of multiple opponents. In our work, we further use the diffusion model as a trajectory planner that allows the reinforcement learning agent to generate diverse plan options in our partially observable pursuit-evasion domain.
II-C RL in Motion Planning
Recent work has demonstrated the success of RL for motion planning problems in dynamic environments [38, 39, 40, 15]. Many papers discuss how to use RL to avoid obstacles by analyzing exteroceptive sensor information such as RGB cameras [40], LIDARs [39, 40], or converted laser data from images [38]. However, none of these previous works assumes an adversarial environment where the “obstacles” are intelligent pursuers, or the map is as large as in our scenario with undetectable parts.
The work [15] also employs a hierarchical structure in goal-conditioned planning; however, their environment does not incorporate adversaries or multiple goals, and their RL policy is not trained from scratch. Dynamic obstacles are considered in [41]; however, these obstacles can only run on a predefined segment with a constant velocity, and the map is known such that A* search can be sufficient to find the optimal path. In our scenario, we assume the cameras are hidden from the evader in a non-discovered terrain map, while the pursuit team can use these cameras to detect and track the evader.
@@ III-A Partially Observable Markov Decision Process
We model adversarial search and tracking as a Partially Observable Markov Decision Process (POMDP). A POMDP for an agent can be defined by a state $s$, private observation $o$, action $a$, and a transition function $\mathcal{T}:s\times a\mapsto s$. At each timestep $t$, an agent receives an observation $o^{t}$, chooses an action $a^{t}\in\mathcal{A}$, and obtains a reward $r^{t}:s\times a_{i}\mapsto\mathbb{R}$. The initial world state is drawn from a prior distribution $\rho$. In our environment (see §III-B), we assume that the evader starts from a random position from a corner region of the map while the pursuit team starts from random positions on the map. It is a partially observable environment for the evader in that 1) cameras are invisible to the evader, 2) the terrain map is not provided a prior, and 3) the evader may only have a limited detection range.
III-B Prisoner Escape and Narco Interdiction
In our evaluation of the proposed algorithm, we employ two open-sourced large-scale pursuit-evasion domains [12] named Prisoner Escape and Narco Interdiction (Figure 2). In both domains, an evader (prisoner or smuggler) is being tracked by a team of search agents in a partially observable environment. In each episode, three hideouts are drawn from all candidate hideouts. These hideouts are only visible to the evader and unknown to the pursuit team, as adversarial teams do not communicate intent. The terminal objective of the evader is to reach any hideout on the map (considered a success) or reach the time limit (considered a failure). Simultaneously, the evader should learn to avoid pursuit team detection. The heterogeneous pursuit team consists of static (cameras) and dynamic agents (search parties, helicopters, search boats, and planes) that collaborate to search and track the evader in both domains.
The maximum speed of the evader is always lower than the pursuit agents, making it more challenging for the evader to plan evasive maneuvers to successfully reach the hideouts undetected. The observation of the evader includes the current timestep, hideout locations, mountain locations, pursuer states, and evader states. The evader can see the pursuer anywhere in the global-view mode and within a field-of-view in the local-view mode, and the camera locations and visibility map are always hidden from the evader. In the local-view mode, evader’s field-of-view, $d_{e}$, is parameterized by the searching agent type, $\beta_{type}$, visibility, $c_{v}$, and searching agent speed, $s_{s}$: $d_{e}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{s}+\eta)$. These domains increase the complexity of the task and more closely match real-world dynamics compared to other pursuit-evasion games [42, 43, 23].
@@ III-C Pursuit Team Behavior
To provide a fair comparison between our evading algorithm and the baselines, we use a fixed heuristic policy for the pursuit team inspired by prior works [44, 45] that achieved good performance in the prisoner escape domain [14]. The search agents can choose to either 1) go to the last detection, 2) search along the predicted path, or 3) search in the vicinity of the last detection. Searching team detection range, $d_{s}$, is parameterized by the agent type, $\beta_{type}$, visibility, $c_{v}$, and evader speed $s_{e}$: $d_{s}=\alpha\cdot(\beta_{type}\cdot c_{v}\cdot s_{e}+\eta)$, which means a higher visibility and evader speed will lead to larger detection range. The pursuer policy is designed to effectively utilize the available information and adapt to different scenarios.
@@ IV Method
The evader’s objective is to reach one of the hideouts on the map and evade detection by the pursuit team, using only partial map knowledge, without any information about camera locations or models for whether the pursuers can detect the evader. RL or diffusion-based motion planning each partially fits our setting: RL can hypothetically learn to evade through trial-and-error. However, RL has difficulties with exploration in the large, adversarial, and multi-goal domains. On the other hand, diffusion-based motion planning can generate diverse plans that reach the goals; however, diffusion models struggle to learn evasive behaviors and avoid the pursuing agents. The reason is that the diffusion model has a relatively long denoising process that is difficult to handle environment dynamics. Additionally, partial observability induces high variance in training.
We propose to leverage the strengths of diffusion and RL policies by using diffusion models to guide RL exploration during training and generate path candidates at inference. We decompose the hierarchy into the following: learning diffusion-based global plans that meet start, terminal, and obstacle constraints from RRT* paths (§IV-A), employing RL local evasive behaviors to avoid detection en-route (§IV-B), and path selection from posteriorly constructed costmap during inference (§IV-C).
@@ 3             $(\mu^{i},\Sigma^{i})\leftarrow s_{\theta}(\tau^{i}),\Sigma_{\theta}(\tau^{i})$
IV-A Diffusion Based Global Plan
In this section, we describe how we leverage diffusion models as a generative motion planner to provide a prior for the RL agent to traverse in large, partially observable, and adversarial environments [46]. We utilize a diffusion model to clone the distribution from RRT* planner and generate waypoints of diverse candidate paths that the agent should follow approximately. This diffusion path planner guides the exploration, increases the sample-time efficiency, and improves the flexibility of the downstream RL algorithm. The underlying insight is to constrain the RL exploration at a promising distribution implicitly encoded by the diffusion model considering the map constraints.
The diffusion path planning should have three key attributes: 1) global constraints are satisfied (start and terminal locations, obstacles, etc.); 2) generated plans are diverse and multimodal and 3) the sampling time of the diffusion model is low. We guide the diffusion sampling procedure using constraint guidance [13] to best satisfy the constraints and use a small number of sampling timesteps to keep the sampling time low. The low sampling time allows us to run the diffusion model during the RL training procedure, which is difficult for traditional search-based planners. Additionally, the parallel sampling reduces the overhead at inference time and provides diverse candidates for the agent to use.
We outline the diffusion training process in Algorithm 1. We first generate paths with RRT* (Line 1) — a sample-based motion planner that yields varied paths even for the same initial and target states, which enables the diffusion model to similarly generate diverse paths. Then we uniformly downsample the path into sparse waypoints (Line 1) and train a diffusion model based on these waypoints (Line 1-1). Generating a path with a few waypoints is sufficient enough to act as the high-level planner as we expect the RL policy to act as the low-level dynamic actor to move between waypoints. Generating fewer waypoints also allows sampling time to be faster, which is crucial for training in an RL loop.
The sampling process is summarized in Algorithm 2 with diffusion path generator, $\pi^{d}$. We apply global constraints at each denoising step (Line 2) to help generate a valid path from start to goal.
@@ IV-A Diffusion Based Global Plan
IV-B RL Based Adversarial Behavior Learning
In addition to our diffusion model considering the global objectives, it is also essential to design an algorithm directly interacting with the environment that can not only accomplish the tasks from the high-level planners but also capture the dynamic patterns in the map to achieve the local adversarial detection avoiding goals. Although evading heuristics exist [12], we believe that a data-driven reinforcement learning algorithm is a more convenient way compared with the hand-crafted policy and can perform better within specific environment distribution.
Therefore, we employ an RL algorithm Soft Actor-Critic (SAC) as the low-level policy with the following aims: 1) approximately follow the waypoints from the diffusion path and 2) adapt the evader velocities to escape from searching agents. We summarize RL training in Algorithm 3 where we augment the observation with the current desired waypoint from the high-level planner (Line 3, 3). The training objective can be formulated as Equation (1) where $\approx$ means a similar low-level policy can be derived from both sides.